OpenStudy (anonymous):

Write an expression that is equivalent to this expression, using the related acute angle. cos (13pi/12)

OpenStudy (anonymous):

I got cos \[\cos \frac{ 11\pi}{ 12 }\]

OpenStudy (anonymous):

i wanna be sure if im right